OpenStudy (anonymous):

find f(g(x)) and g(f(x)) andetermine whether the pair of functions f and g are inverses of each other f(x)=2x and g(x)=x/2

OpenStudy (anonymous):

f[g(x)]=f[x/2]=2(x/2)=x and g[f(x)]=g[2x]=(2x)/2=x Thus f and g are inverse functions